OpenStudy (anonymous):

Given f(x)=√x-1 find F (1), F (2), F(5) Please help

OpenStudy (anonymous):

replace \(x\) by \(1\) and compute \[f(1)=\sqrt{1-1}=\sqrt{0}=0\]